OpenStudy (anonymous):

how long would it take the voyager, traveling at warp 8, to cross the entire radius of the milky way?

OpenStudy (anonymous):

the radius of the milky way is 2.94 x 10^17 miles

OpenStudy (anonymous):

Warp 8 = 1024c = 186,000*1024 = 190,464,000 miles per second. Speed = Distance/time 190,464,000 = 2.94 x 10^17/x You should be able to take it from there. x = time in seconds.

OpenStudy (anonymous):

where did you get the 186,000?

OpenStudy (anonymous):

It's the speed of light in miles/sec

OpenStudy (anonymous):

It's 186,282.397 miles/second to be exact, but mostly it's rounded to 186,000 for ease of use.

OpenStudy (anonymous):

That's meters per second. It's the same speed, just in a different unit. If you converted one number to the other (i.e. meters to miles or miles to meters) you would find them equivalent.

OpenStudy (anonymous):

It would also work just as well. Instead of substituting C = 186,000 miles/second, you would substitute C = 299,792, 458 meters/second. You'd still get the same time (as long as you remember to convert your galaxy radius distance to meters as well. You always must work in the same unit!).

OpenStudy (anonymous):

I got 1.019702238*10^25, is that correct?

OpenStudy (anonymous):

It depends on which calculation you did. Did you use meters or miles for the distance?

OpenStudy (anonymous):

meters

OpenStudy (anonymous):

it also says on the worksheet that warp 8 = 8*the speed of light

OpenStudy (anonymous):

Well, in that case, replace the "1024" with "8" and then tell your teacher that Warp 8 in real Star Trek isn't 8x the speed of light, it's 1024x :P

OpenStudy (anonymous):

so would i do 186,282.397 miles/seconds * 8? and then divide by 2.94*10^17?

OpenStudy (anonymous):

okay and will do:)

OpenStudy (anonymous):

Speed = Distance/time. To find time, we do Time = Distance/Speed

OpenStudy (anonymous):

So if you're going to work in miles, you would do 186,282.397 * 8 to get your speed at Warp 8 in miles/second. Your distance is a given. Just plug it in after that. If you're going to work in meters, it would be 299,792,458 * 8 to get your speed at Warp 8 in meters/second. Then you would have to convert the distance from miles to meters (hint: there are 1609.34 meters in a mile), then plug in. Doing it either way should give you the same answer.

OpenStudy (anonymous):

so it would be 8*186,282,397=1790259176m/s, which is the distance and dividing it by time would mean dividing 1790259176m/s by 2.94*10^17=6.089316925*10^25?

OpenStudy (anonymous):

The 8*186,282.397 (be careful, it's not one hundred and eight six million, it's one hundred and eighty six thousand. There is only one comma, the other mark is a decimal point) is your speed, not distance. The distance is the given, the radius of the Milky Way Galaxy. Time is your unknown, that's what you're looking for.

OpenStudy (anonymous):

never mind i figured it out! thanks for your help!!!!!!

OpenStudy (anonymous):

just to make sure the answer is 5.068908762*10^22 seconds?

OpenStudy (anonymous):

No, that's not what I got. If you write out your calculations, I can show you where you're going wrong.

OpenStudy (anonymous):

okay, 8.186,282.397*8=1490259.176m/s t=d/s t=2.94*10^17miles/1490259.176m/s=1.972811204*10^11

OpenStudy (anonymous):

Yep, that's what I got. That is the correct answer for how many seconds it would take.

OpenStudy (anonymous):

so if the question says how long would it take the voyager, traveling at warp 8, to cross the entire radius of the milky way? 1.972811204*10^11 this would be a correct answer?

OpenStudy (anonymous):

Don't forget your units! 1.972811204*10^11 SECONDS. Yes, that is the correct answer.

OpenStudy (anonymous):

thats what i meant but yeah thank you once again for your help!!!!!!!!!!!

OpenStudy (anonymous):

No problem. Anytime.